OpenStudy (anonymous):

If A is a subset of B, prove that P(B|A)=1

OpenStudy (anonymous):

probability

OpenStudy (ash2326):

@SKMC It's given that A is a subset of B so all the elements of A are also in B we need to find the probability of occurrence of B given that A has occurred, now A has already occurred and also A is a subset of B, so B would have occurred too, so probability will always be 1

OpenStudy (ash2326):

@SKMC did you understand?

OpenStudy (ash2326):

I can explain you using an example, do you want me to do so?

OpenStudy (anonymous):

I understand the concept, I am just trying to prove it in the correct notation So far I have the following...

OpenStudy (anonymous):

\[P(B|A) = P(AB)P(B)/P(A)\]

OpenStudy (anonymous):

So i need to get to \[P(A) + P(B) =1\]

OpenStudy (ash2326):

Oh, your approach is right but \[P(B|A)=P(A\cap B)/P(B)\] What is \(P(A\cap B )\)?

OpenStudy (ash2326):

@SKMC A is a subset of B, so what's \(P(A\cap B)\)?

OpenStudy (anonymous):

A & B simultaneously

OpenStudy (ash2326):

That's right but what it boils down to?

OpenStudy (anonymous):

You may like to state what \( A\) and \( B \) is in your question.

OpenStudy (ash2326):

@FoolForMath A= red balls B= red balls+ black balls C= blue balls + pink balls Given that a red ball is drawn (which means A has occurred) , what's the probability that a ball from B is drawn. A is a subset of B so B has also occurred so P(B|A)=1

OpenStudy (anonymous):

Unless I am going mad, \[ P(B|A) \neq P(A\cap B)/P(B) \]

OpenStudy (ash2326):

A is a subset of B \[P(B/A)= P (A\cap B)/P(A)\] now \[P(A\cap B)= P(A)\] \[P(B/A)= P (A)/P(A)=1\]

OpenStudy (anonymous):

Yes, If \( A\subset B \implies P(A\cap B) = P(A)\) that's all we are required in this problem.

OpenStudy (anonymous):

Many Thanks for your help. I knew I was missing something that should have been obvious - really appreciate your help.

OpenStudy (anonymous):

Glad to help :)